OpenStudy (anonymous):

solve for x: 4^x - 3*2^(x+1) = 16

OpenStudy (anonymous):

\[4^{x} - 3 * 2^{x+1} =16\]

OpenStudy (anonymous):

Hint: take log on both sides

OpenStudy (anonymous):

can you solve for me. i guarantee a medal

OpenStudy (anonymous):

\[\log(4^{x}) -3.\log(2^{x+1}) = \log(16)\] \[x.\log4 - 3 (x+1) \log2 = \log 16\]

OpenStudy (john_es):

log functin won't work here.

OpenStudy (anonymous):